Could you please help me solve these systems by using the addition method and indicate whether each system is independent, inconsistent, or dependent.
x+3(y-1)=11
2(x-y)+8y-28
Thank you!!!

Answer by stanbon(48545)

solve these systems by using the addition method and indicate whether each system is independent, inconsistent, or dependent.
x+3(y-1)=11
2(x-y)+8y-28
-----------
Rewrite as:
x+3y=14
2x+6y=28
-------------
Multiply the 1st equation be 2 and subtract the results from the 2nd equation:
You get:
0 = 0
This happened because the two equation are the same.
So the system is dependent.
The solution set for the system is the line y= (-1/3)x+(14/3)
===============
Cheers,
Stan H.